Demonstrate that 2+6i is a solution to the equation x^2-40x+40=0
Answer by venugopalramana(3286)

Demonstrate that 2+6i is a solution to the equation x^2-40x+40=0
SUBSTITUTING FOR X ..
(2+6i)^2-40(2+6i)+40
=4+36i^2+2*2*6i-80-240i+40
=4-36+24i-80-240i+40=
=NO...IT IS NOT A SOLUTION AS IT IS NOT ZERO